OpenStudy (anonymous):

In your class 11 out of 25 students wear glasses.What percent if the students have wear glasses? Work it into a proportion ASAP

jimthompson5910 (jim_thompson5910):

11/25 = x/100 11*100 = 25x 25x =1100 x = ??